# Causal reasoning in difference graphs

### v1

@@ 3 Identification of total effects using difference graphs in a nonparametric setting
Consider a difference graph $\mathcal{D}$ compatible with two different SCMs. Under Assumptions 1, 2 and 3, $\Pr(y\mid\text{do}(x))$ is identifiable in $\mathcal{D}$ by a common back-door if and only if:
@@
    $X\in Ancestors(Y,\mathcal{D})$ and $\forall W\in\mathbb{V}\backslash\{X,Y\}$, $X\in Ancestors(W,\mathcal{D})$ or $W\in Ancestors(X,\mathcal{D})$.
Furthermore, if Condition A.1 is satisfied then $\Pr(y\mid\text{do}(x))=\Pr(y)$ and if Condition A.2 is satisfied then $\Pr(y\mid\text{do}(x))=\sum_{\mathbb{w}^{anc}}\Pr(y\mid x,\mathbb{w}^{anc})\Pr(\mathbb{w}^{anc})$, where $\mathbb{W}^{anc}=Ancestors(X,\mathcal{D})\backslash\{X\}$.
@@
In particular, when there is no directed path between the variables of interest—either from $X$ to $Y$ or from $Y$ to $X$—as seen in the difference graph in Figure 1 (c), the absence of such a connection creates ambiguity in the directionality of the total effect. This is illustrated by the pair of causal DAGs in (a) and (b), where the association between $X$ and $Y$ reflects the total effect of interest, and by the pair in (d) and (e), where the association between $X$ and $Y$ does not correspond to the total effect being sought. Therefore, the total effect cannot be identified using a common back-door. This is an essential point, as it highlights that causal reasoning in difference graphs can appear counterintuitive, with nonexistent directed paths in the difference graph potentially leading to ambiguous causal interpretations. The following lemma builds on this observation and in addition it shows that Condition A.1 of the theorem is sufficient.
@@
Condition A.2 of Theorem 1 can be applied to identify $\Pr(y\mid\text{do}(x))$ in the difference graph shown in Figure 1 (h), as $W_{1}\in Ancestors(X,\mathcal{D})$ and $X\in Ancestors(W_{2},\mathcal{D})$. This implies that for any pair of causal DAGs compatible with the difference graph—such as the pairs depicted in Figures 1 (f) and (g) or Figures 1 (i) and (j)—the set $\{W_{1}\}$ satisfies the back-door criterion. The following lemma illustrates that Condition A.2 is indeed sufficient for identifiability.
@@
Consider any pair of causal DAGs ($\mathcal{G}^{1}$, $\mathcal{G}^{2}$) that define $\mathcal{D}$. Let $\mathbb{W}^{anc}=Ancestors(X,\mathcal{D})\backslash\{X,Y\}$ and let $\mathbb{W}^{desc}=Descendants(X,\mathcal{D})\backslash\{X,Y\}$ For any $W\in\mathbb{W}^{anc}$, $W$ cannot be in $\mathbb{W}^{desc}$, otherwise $\mathcal{D}$ would not be acyclic and Assumption 3 would not be satisfied. By the lemma, for any $W\in\mathbb{V}\backslash\{X,Y\}$ if $W\notin\mathbb{W}^{anc}$ then $W\in\mathbb{W}^{desc}$. It follows that $\forall W\in\mathbb{W}^{anc}$, $W$ has to be an ancestor of $X$ in at least one of the two causal DAGs $\mathcal{G}^{1}$ and $\mathcal{G}^{2}$ and $W$ is not a descendent of $X$ in any of the two causal DAGs. Thus $\mathbb{W}^{anc}$ has to block all paths from $X$ to $Y$ which contains an arrow into $X$ and at the same time $\mathbb{W}^{anc}$ cannot contain any descendant of $X$. Therefore, the set $\mathbb{W}^{anc}$ satisfies the backdoor criterion [Pearl, 1995] in $\mathcal{G}^{1}$ and in $\mathcal{G}^{2}$ and by Pearl [1995, Theorem 1] the total effect $\Pr(y\mid\text{do}(x))$ in $\mathcal{M}^{1}$ and in $\mathcal{M}^{2}$ is given by $\sum_{\mathbb{w}^{anc}}\Pr(y\mid x,\mathbb{w}^{anc})\Pr(\mathbb{w}^{anc})$. ∎
@@
Under Assumptions 1, 2 and 3, if there exists a directed path from $X$ to $Y$ and there exists a vertex $W\in\mathbb{V}\backslash\{X,Y\}$ for which there is no directed path from $X$ to $W$ or from $W$ to $X$ then $\Pr(y\mid\text{do}(x))$ is not identifiable in $\mathcal{D}$ by a common back-door.
@@
Suppose there is a directed path from $X$ to $Y$. Now, consider a vertex $W$ for which there is no directed path from or to $X$ in $\mathcal{D}$. It follows that in any pair of causal DAGs $\mathcal{G}^{1}$ and $\mathcal{G}^{2}$, $W$ can be either independent of $X$ or a descendant of $X$ or a parent of $X$ or an ancestor, but not a parent, of $X$. In the first and fourth cases, $W$ can be included in a set that satisfies the back-door criterion in all causal DAGs compatible with $\mathcal{D}$. In the second case, $W$ cannot be included in any set that satisfies the back-door criterion in every causal DAG compatible with $\mathcal{D}$, since $W$ is a descendant of $X$. In the third case, $W$ cannot be a descendant of $Y$; otherwise, a cycle would form involving $X$, $W$, and $Y$. Therefore, in at least one causal DAG compatible with $\mathcal{D}$, $W$ must be a parent of $Y$, meaning $W$ must be included in any set that satisfies the back-door criterion in every causal DAG compatible with $\mathcal{D}$. This final case contradicts the second case, making it impossible to find a set that satisfies the back-door criterion in all compatible causal DAGs. ∎
The remaining part of this section investigates the case where the two causal DAGs induced by the two SCMs that define the difference graph do not share the same topological oredering (i.e., Assumption 3 is not satisfied). In this case the difference graph can contain cycles.
@@
Consider a difference graph $\mathcal{D}$ compatible with two different SCMs. Under Assumptions 1 and 2, $\Pr(y\mid\text{do}(x))$ is identifiable in $\mathcal{D}$ by a common back-door if and only if
@@
    Condition A.1 is satisfied and $X\not\in Ancestors(Y,\mathcal{D})$; or
@@
    Condition A.2 is satisfied and $Ancestors(X,\mathcal{D})\cap Descendants(X,\mathcal{D})=\{X\}$.
Furthermore, if Condition B.1 is satisfied then $\Pr(y\mid\text{do}(x))=\Pr(y)$ and if Condition B.2 is satisfied then $\Pr(y\mid\text{do}(x))=\sum_{\mathbb{w}^{anc}}\Pr(y\mid x,\mathbb{w}^{anc})\Pr(\mathbb{w}^{anc})$, where $\mathbb{W}^{anc}=Ancestors(X,\mathcal{D})\backslash\{X\}$.
@@
The proof is very similar to the proof of Theorem 1 with sime minor difference. First, notice that replacing Assumption 3 with the condition that there is no directed path from $X$ to $Y$ in the proof of Lemma 3.1 would maintain correctness. Second notice that replacing Assumption 3 with the condition that there is no cycle containing $X$ and another vertex in the proof of Lemma 3.2 would maintain correctness. Finally, to extend Lemma 3.3, we add the condition that the total effect is not identifiable by a common back-door if a cycle exists on $X$. Combining this with the second part of Lemma 3.1 establishes that the conditions are necessary. This can be easily shown because if there is a cycle in $\mathcal{D}$ involving $X$ and another vertex $W$ (where $W\neq Y$), then there must exist a causal DAG in which $W$ is a parent of $X$, as well as another causal DAG in which $W$ is a descendant of $X$. In both of these graphs, $W$ can also be a parent of $Y$. This implies that in the first DAG, $W$ must be included in any set satisfying the back-door criterion, while in the second DAG, $W$ cannot be included in such a set. Therefore, no set can satisfy the back-door criterion for all causal DAGs compatible with $\mathcal{D}$. Moreover, if the cycle only involves $X$ and $Y$, the total effect is also not identifiable for the same reason the total effect is not identifiable when there is no directed path from $X$ to $Y$ or from $Y$ to $X$. ∎
@@ 4 Mediation analysis using difference graphs in a linear setting
Consider a difference graph $\mathcal{D}$ compatible with two different linear SCMs. Under Assumptions 1, 2 and 3, $\alpha_{x,y}$ is identifiable in $\mathcal{D}$ by a common single-door if and only if:
@@
    $X\in Ancestors(Y,\mathcal{D})$ and $\forall W\in\mathbb{V}\backslash\{X,Y\}$, $Y\in Ancestors(W,\mathcal{D})$ or $W\in Ancestors(Y,\mathcal{D})$.
Furthermore, if Condition C.1 is satisfied then $\alpha_{x,y}=0$ and if Condition C.2 is satisfied then $\alpha_{x,y}=r_{YX.\mathbb{W}^{anc}}$ where $r_{YX.\mathbb{W}^{anc}}$ is the regression coefficient of $X$ on $Y$ after $\mathbb{W}^{anc}$ is partialled out.
@@
Theorem 3 establishes that the direct effect $\alpha_{x,y}$ is identifiable in the difference graph presented in Figure 1 (m), but it is not identifiable using common single-door in the difference graphs shown in Figures 1 (c) and (h). The same reasoning applied in the previous section for the non-identifiability of the total effect in Figure 1 (c) also applies here for the direct effect. For the case of Figure 1 (m), it becomes clear why the direct effect is not identifiable when considering the two pairs of causal DAGs compatible with the difference graph, shown in Figures (f) and (g), and Figures (i) and (j). In the first pair, $\{W_{1},W_{2}\}$ is the only set that satisfies the single-door criterion. However, in the second pair, $W_{2}$ cannot be included in any set that satisfies the single-door criterion. We now move on to present the lemmas required for the proof of the theorem, noting that their proofs follow the same reasoning as the lemmas introduced in the previous section.
@@
Consider any pair of causal DAGs ($\mathcal{G}^{1}$, $\mathcal{G}^{2}$) that define $\mathcal{D}$. Let $\mathbb{W}^{anc}=Ancestors(Y,\mathcal{D})\backslash\{X,Y\}$ and let $\mathbb{W}^{desc}=Descendants(Y,\mathcal{D})\backslash\{X,Y\}$ For any $W\in\mathbb{W}^{anc}$, $W$ cannot be in $\mathbb{W}^{desc}$, otherwise $\mathcal{D}$ would not be acyclic and Assumption 3 would not be satisfied. By the lemma, for any $W\in\mathbb{V}\backslash\{X,Y\}$ if $W\notin\mathbb{W}^{anc}$ then $W\in\mathbb{W}^{desc}$. It follows that $\forall W\in\mathbb{W}^{anc}$, $W$ has to be an ancestor of $Y$ in at least one of the two causal DAGs $\mathcal{G}^{1}$ and $\mathcal{G}^{2}$ and $W$ is not a descendent of $Y$ in any of the two causal DAGs. Thus $\mathbb{W}^{anc}$ has to block all paths between $X$ and $Y$ and at the same time $\mathbb{W}^{anc}$ cannot contain any descendant of $Y$. Therefore, the set $\mathbb{W}^{anc}$ satisfies the single-door criterion [Spirtes et al., 1998, Pearl, 1998] in $\mathcal{G}^{1}$ and in $\mathcal{G}^{2}$ and by Pearl [1998, Theorem 6] the direct effect $\alpha_{x,y}$ in $\mathcal{M}^{1}$ and in $\mathcal{M}^{2}$ is given by $r_{Y,X.\mathbb{W}^{anc}}$. ∎
@@
Suppose there is a directed path from $X$ to $Y$. Now, consider a vertex $W$ for which there is no directed path from or to $Y$ in $\mathcal{D}$. It follows that in any pair of causal DAGs $\mathcal{G}^{1}$ and $\mathcal{G}^{2}$, $W$ can be either independent of $Y$ or a descendant of $Y$ or a parent of $Y$ or an ancestor, but not a parent, of $Y$. In the first and fourth cases, $W$ can be included in a set that satisfies the single-door criterion in all causal DAGs compatible with $\mathcal{D}$. In the second case, $W$ cannot be included in any set that satisfies the single-door criterion in every causal DAG compatible with $\mathcal{D}$, since $W$ is a descendant of $Y$. In the third case, it is possible to imagine that thre exists a causal DAG compatible with $\mathcal{D}$ where $W$ is either a parent of a child of $X$. Therefore, in this causal DAG, $W$ must be included in any set that satisfies the single-door criterion in every causal DAG compatible with $\mathcal{D}$. This final case contradicts the second case, making it impossible to find a set that satisfies the single-door criterion in all compatible causal DAGs. ∎
@@
Consider a difference graph $\mathcal{D}$ compatible with two different linear SCMs. Under Assumptions 1 and 2, $\alpha_{x,y}$ is identifiable in $\mathcal{D}$ by a common single-door if and only if:
@@
    Condition C.1 is satisfied and $X\not\in Ancestors(Y,\mathcal{D})$; or
@@
    Condition C.2 is satisfied and $Ancestors(Y,\mathcal{D})\cap Descendants(Y,\mathcal{D})=\{Y\}$.
Furthermore, if Condition D.1 is satisfied then $\alpha_{x,y}=0$ and if Condition D.2 is satisfied then $\alpha_{x,y}=r_{YX.\mathbb{W}^{anc}}$ where $r_{YX.\mathbb{W}^{anc}}$ is the regression coefficient of $X$ on $Y$ after $\mathbb{W}^{anc}$ is partialled out.
@@
The proof closely follows the structure of the proof for Theorem 3, with a few key modifications. First, note that replacing Assumption 3 with the condition that there is no directed path from $X$ to $Y$ in the proof of Lemma 4.1 preserves its correctness. Second, replacing Assumption 3 with the condition that no cycle contains both $Y$ and another vertex in the proof of Lemma 4.2 also maintains its validity. To extend Lemma 4.3, we introduce the condition that the total effect cannot be identified using a common single-door criterion if there is a cycle involving $Y$. Combining this new condition with the second part of Lemma 4.1 demonstrates that these conditions of the Theorem are necessary. This can be shown as follows: if a cycle exists in $\mathcal{D}$ involving $Y$ and another vertex $W$ (where $W\neq X$), then there must be at least one causal DAG where $W$ is a parent of $Y$ and another causal DAG where $W$ is a descendant of $Y$. In both of these DAGs, $W$ can also act as either a parent or descendant of $X$. This implies that in the first DAG, $W$ must be included in any set that satisfies the single-door criterion, while in the second DAG, $W$ cannot be included in such a set. As a result, no single set can satisfy the single-door criterion across all causal DAGs compatible with $\mathcal{D}$. Furthermore, if the cycle only involves $X$ and $Y$, the direct effect is also not identifiable for the same reason it is not identifiable when there is no directed path between $X$ and $Y$. ∎
@@
5 Related works

### v2

@@ 3 Identification of total effects using difference graphs in a nonparametric setting
Consider a difference graph $\mathcal{D}$ compatible with two different SCMs. Under Assumptions 1, 2 and 3, $\Pr(y\mid\text{do}(x))$ is identifiable in $\mathcal{D}$ trivially or by a common back-door iff:
@@
    $X\in An(Y,\mathcal{D})$ and $\forall W\in\mathbb{V}\backslash\{X,Y\}$, $X\in An(W,\mathcal{D})$ or $W\in An(X,\mathcal{D})$.
Furthermore, if Condition A.1 is satisfied then $\Pr(y\mid\text{do}(x))=\Pr(y)$ and if Condition A.2 is satisfied then $\Pr(y\mid\text{do}(x))=\sum_{\mathbb{w}^{an}}\Pr(y\mid x,\mathbb{w}^{an})\Pr(\mathbb{w}^{an})$, where $\mathbb{W}^{an}=An(X,\mathcal{D})\backslash\{X\}$.
@@
In particular, when there is no directed path between the variables of interest—either from $X$ to $Y$ or from $Y$ to $X$—as seen in the difference graph in Figure 1 (c), the absence of such a connection creates ambiguity in the directionality of the total effect. This is illustrated by the pair of causal DAGs in (a) and (b), where the association between $X$ and $Y$ reflects the total effect of interest, and by the pair in (d) and (e), where the association between $X$ and $Y$ does not correspond to the total effect being sought. Therefore, the total effect cannot be identified using a common back-door. This is an essential point, as it highlights that causal reasoning in difference graphs is counterintuitive, with nonexistent directed paths in the difference graph potentially leading to ambiguous causal interpretations. The following lemma builds on this observation and in addition it shows that Condition A.1 of the theorem is sufficient.
@@
Condition A.2 of Theorem 1 can be applied to identify $\Pr(y\mid\text{do}(x))$ in the difference graph shown in Figure 1 (h), as $W_{1}\in An(X,\mathcal{D})$ and $X\in An(W_{2},\mathcal{D})$. This implies that for any pair of causal DAGs compatible with the difference graph—such as the pairs depicted in Figures 1 (f) and (g) or Figures 1 (i) and (j)—the set $\{W_{1}\}$ satisfies the back-door criterion. The following lemma illustrates that Condition A.2 is indeed sufficient for identifiability.
@@
Consider any pair of causal DAGs ($\mathcal{G}^{1}$, $\mathcal{G}^{2}$) that define $\mathcal{D}$. Let $\mathbb{W}^{an}=An(X,\mathcal{D})\backslash\{X,Y\}$ and let $\mathbb{W}^{de}=De(X,\mathcal{D})\backslash\{X,Y\}$. For any $W\in\mathbb{W}^{an}$, $W$ cannot be in $\mathbb{W}^{de}$, otherwise $\mathcal{D}$ would not be acyclic and Assumption 3 would not be satisfied. By the lemma, for any $W\in\mathbb{V}\backslash\{X,Y\}$ if $W\notin\mathbb{W}^{an}$ then $W\in\mathbb{W}^{de}$. It follows that $\forall W\in\mathbb{W}^{an}$, $W$ has to be an ancestor of $X$ in at least one of the two causal DAGs $\mathcal{G}^{1}$ and $\mathcal{G}^{2}$ and $W$ is not a descendent of $X$ in any of the two causal DAGs. Thus $\mathbb{W}^{an}$ has to block all paths from $X$ to $Y$ which contains an arrow into $X$ and at the same time $\mathbb{W}^{an}$ cannot contain any descendant of $X$. Therefore, $\mathbb{W}^{an}$ satisfies the back-door criterion [Pearl, 1995] in $\mathcal{G}^{1}$ and in $\mathcal{G}^{2}$ and by Pearl [1995, Theorem 1] the total effect $\Pr(y\mid\text{do}(x))$ in $\mathcal{M}^{1}$ and in $\mathcal{M}^{2}$ is given by $\sum_{\mathbb{w}^{an}}\Pr(y\mid x,\mathbb{w}^{an})\Pr(\mathbb{w}^{an})$. ∎
@@
Under Assumptions 1, 2 and 3, if there exists a directed path from $X$ to $Y$ and there exists a vertex $W\in\mathbb{V}\backslash\{X,Y\}$ for which there is no directed path from $X$ to $W$ or from $W$ to $X$ in $\mathcal{D}$ then $\Pr(y\mid\text{do}(x))$ is not identifiable in $\mathcal{D}$ by a common back-door.
@@
Suppose there is a directed path from $X$ to $Y$ in $\mathcal{D}$. Now, consider a vertex $W$ for which there is no directed path from or to $X$ in $\mathcal{D}$. It follows that in any pair of causal DAGs $\mathcal{G}^{1}$ and $\mathcal{G}^{2}$, $W$ can be either independent of $X$ or a descendant of $X$ or a parent of $X$ or an ancestor, but not a parent, of $X$. Let us focus on the second and third cases. In the second case, $W$ cannot be included in any set that satisfies the back-door criterion in every causal DAG compatible with $\mathcal{D}$, since $W$ is a descendant of $X$. In the third case, $W$ cannot be a descendant of $Y$; otherwise, a cycle would form involving $X$, $W$, and $Y$. Therefore, in at least one causal DAG compatible with $\mathcal{D}$, $W$ must be a parent of $Y$, meaning $W$ must be included in any set that satisfies the back-door criterion in every causal DAG compatible with $\mathcal{D}$. This third case contradicts the second case, making it impossible to find a set that satisfies the back-door criterion in all compatible causal DAGs.
The remaining part of this section investigates the case where the two causal DAGs induced by the two SCMs that define the difference graph do not share the same topological ordering (i.e., Assumption 3 is not satisfied). In this case the difference graph can contain cycles.
@@
Consider a difference graph $\mathcal{D}$ compatible with two different SCMs. Under Assumptions 1 and 2, $\Pr(y\mid\text{do}(x))$ is identifiable in $\mathcal{D}$ trivially or by a common back-door iff
@@
    Condition A.1 is satisfied and $X\not\in An(Y,\mathcal{D})$; or
@@
    Condition A.2 is satisfied and $An(X,\mathcal{D})\cap De(X,\mathcal{D})=\{X\}$.
Furthermore, if Condition B.1 is satisfied then $\Pr(y\mid\text{do}(x))=\Pr(y)$ and if Condition B.2 is satisfied then $\Pr(y\mid\text{do}(x))=\sum_{\mathbb{w}^{an}}\Pr(y\mid x,\mathbb{w}^{an})\Pr(\mathbb{w}^{an})$, where $\mathbb{W}^{an}=An(X,\mathcal{D})\backslash\{X\}$.
@@
The proof is very similar to the proof of Theorem 1 with sime minor difference. First, notice that replacing Assumption 3 with the condition that there is no directed path from $X$ to $Y$ in the proof of Lemma 3.1 would maintain correctness. Second notice that replacing Assumption 3 with the condition that there is no cycle containing $X$ and another vertex in the proof of Lemma 3.2 would maintain correctness. Finally, to extend Lemma 3.3, I add the condition that the total effect is not identifiable by a common back-door if a cycle exists on $X$. Combining this with the second part of Lemma 3.1 establishes that the conditions are necessary. This can be easily shown because if there is a cycle in $\mathcal{D}$ involving $X$ and another vertex $W$ (where $W\neq Y$), then there must exist a causal DAG in which $W$ is a parent of $X$, as well as another causal DAG in which $W$ is a descendant of $X$. In both of these graphs, $W$ can also be a parent of $Y$. This implies that in the first DAG, $W$ must be included in any set satisfying the back-door criterion, while in the second DAG, $W$ cannot be included in such a set. Therefore, no set can satisfy the back-door criterion for all causal DAGs compatible with $\mathcal{D}$. Moreover, if the cycle only involves $X$ and $Y$, the total effect is also not identifiable for the same reason the total effect is not identifiable when there is no directed path from $X$ to $Y$ or from $Y$ to $X$. ∎
@@
Small simulation study for total causal changes.
The adjustment set provided by Theorems 1 and 2 is not the only possible adjustment set that could be selected using the back-door criterion if the true pair of DAGs were known. However, as shown in Figure 3 (a), in practice, the set used in Theorems 1 and 2 performs as well as the set consisting of the parents of the exposure and outperforms the minimal set derived from the back-door criterion.
@@ 4 Mediation analysis using difference graphs in a linear setting
Consider a difference graph $\mathcal{D}$ compatible with two different linear SCMs. Under Assumptions 1, 2 and 3, $\alpha_{x,y}$ is identifiable in $\mathcal{D}$ trivially or by a common single-door iff:
@@
    $X\in An(Y,\mathcal{D})$ and $\forall W\in\mathbb{V}\backslash\{X,Y\}$, $Y\in An(W,\mathcal{D})$ or $W\in An(Y,\mathcal{D})$.
Furthermore, if Condition C.1 is satisfied then $\alpha_{x,y}=0$ and if Condition C.2 is satisfied then $\alpha_{x,y}=r_{YX.\mathbb{W}^{an}}$ where $r_{YX.\mathbb{W}^{an}}$ is the regression coefficient of $X$ on $Y$ after $\mathbb{W}^{an}$ is partialled out.
@@
Theorem 3 establishes that the direct effect $\alpha_{x,y}$ is identifiable by a common single-door in the difference graph presented in Figure 1 (m), but it is not identifiable by a common single-door in the difference graphs shown in Figures 1 (c) and (h). The same reasoning applied in the previous section for the non-identifiability of the total effect in Figure 1 (c) also applies here for the direct effect. For the case of Figure 1 (m), it becomes clear why the direct effect is not identifiable when considering the two pairs of causal DAGs compatible with the difference graph, shown in Figures (f) and (g), and Figures (i) and (j). In the first pair, $\{W_{1},W_{2}\}$ is the only set that satisfies the single-door criterion. However, in the second pair, $W_{2}$ cannot be included in any set that satisfies the single-door criterion. I now move on to present the lemmas required for the proof of the theorem, noting that their proofs follow the same reasoning as the lemmas introduced in the previous section.
@@
Consider any pair of causal DAGs ($\mathcal{G}^{1}$, $\mathcal{G}^{2}$) that define $\mathcal{D}$. Let $\mathbb{W}^{an}=An(Y,\mathcal{D})\backslash\{X,Y\}$ and let $\mathbb{W}^{de}=De(Y,\mathcal{D})\backslash\{X,Y\}$ For any $W\in\mathbb{W}^{an}$, $W$ cannot be in $\mathbb{W}^{de}$, otherwise $\mathcal{D}$ would not be acyclic and Assumption 3 would not be satisfied. By the lemma, for any $W\in\mathbb{V}\backslash\{X,Y\}$ if $W\notin\mathbb{W}^{an}$ then $W\in\mathbb{W}^{de}$. It follows that $\forall W\in\mathbb{W}^{an}$, $W$ has to be an ancestor of $Y$ in at least one of the two causal DAGs $\mathcal{G}^{1}$ and $\mathcal{G}^{2}$ and $W$ is not a descendent of $Y$ in any of the two causal DAGs. Thus $\mathbb{W}^{an}$ has to block all paths between $X$ and $Y$ and at the same time $\mathbb{W}^{an}$ cannot contain any descendant of $Y$. Therefore, $\mathbb{W}^{an}$ satisfies the single-door criterion [Spirtes et al., 1998, Pearl, 1998] in $\mathcal{G}^{1}$ and in $\mathcal{G}^{2}$ and by Pearl [1998, Theorem 6] the direct effect $\alpha_{x,y}$ in $\mathcal{M}^{1}$ and in $\mathcal{M}^{2}$ is given by $r_{Y,X.\mathbb{W}^{an}}$. ∎
@@
Suppose there is a directed path from $X$ to $Y$. Now, consider a vertex $W$ for which there is no directed path from or to $Y$ in $\mathcal{D}$. It follows that in any pair of causal DAGs $\mathcal{G}^{1}$ and $\mathcal{G}^{2}$, $W$ can be either independent of $Y$ or a descendant of $Y$ or a parent of $Y$ or an ancestor, but not a parent, of $Y$. Let us focus on the second and third cases. In the second case, $W$ cannot be included in any set that satisfies the single-door criterion in every causal DAG compatible with $\mathcal{D}$, since $W$ is a descendant of $Y$. In the third case, it is possible to imagine that thre exists a causal DAG compatible with $\mathcal{D}$ where $W$ is either a parent of a child of $X$. Therefore, in this causal DAG, $W$ must be included in any set that satisfies the single-door criterion in every causal DAG compatible with $\mathcal{D}$. This third case contradicts the second case, making it impossible to find a set that satisfies the single-door criterion in all compatible causal DAGs.
@@
Consider a difference graph $\mathcal{D}$ compatible with two different linear SCMs. Under Assumptions 1 and 2, $\alpha_{x,y}$ is identifiable in $\mathcal{D}$ trivially or by a common single-door iff:
@@
    Condition C.1 is satisfied and $X\not\in An(Y,\mathcal{D})$; or
@@
    Condition C.2 is satisfied and $An(Y,\mathcal{D})\cap De(Y,\mathcal{D})=\{Y\}$.
Furthermore, if Condition D.1 is satisfied then $\alpha_{x,y}=0$ and if Condition D.2 is satisfied then $\alpha_{x,y}=r_{YX.\mathbb{W}^{an}}$ where $r_{YX.\mathbb{W}^{an}}$ is the regression coefficient of $X$ on $Y$ after $\mathbb{W}^{an}$ is partialled out.
@@
The proof closely follows the structure of the proof for Theorem 3, with a few key modifications. First, note that replacing Assumption 3 with the condition that there is no directed path from $X$ to $Y$ in the proof of Lemma 4.1 preserves its correctness. Second, replacing Assumption 3 with the condition that no cycle contains both $Y$ and another vertex in the proof of Lemma 4.2 also maintains its validity. To extend Lemma 4.3, I introduce the condition that the direct effect cannot be identified using a common single-door criterion if there is a cycle involving $Y$. Combining this new condition with the second part of Lemma 4.1 demonstrates that these conditions of the Theorem are necessary. This can be shown as follows: if a cycle exists in $\mathcal{D}$ involving $Y$ and another vertex $W$ (where $W\neq X$), then there must be at least one causal DAG where $W$ is a parent of $Y$ and another causal DAG where $W$ is a descendant of $Y$. In both of these DAGs, $W$ can also act as either a parent or descendant of $X$. This implies that in the first DAG, $W$ must be included in any set that satisfies the single-door criterion, while in the second DAG, $W$ cannot be included in such a set. As a result, no single set can satisfy the single-door criterion across all causal DAGs compatible with $\mathcal{D}$. Furthermore, if the cycle only involves $X$ and $Y$, the direct effect is also not identifiable for the same reason it is not identifiable when there is no directed path between $X$ and $Y$. ∎
@@
Small simulation study for direct causal changes.
The adjustment set provided by Theorems 3 and 4 is not the only possible adjustment set that could be selected using the single-door criterion if the true pair of DAGs were known. So similarly, to Section 3, in Figure 3 (b), the set used in Theorems 3 and 4 is compared with the set consisting of the parents of the outcome and the minimal set derived from the single-door criterion. In this case, the set used in Theorems 3 and 4 performs slightly less well than the set of parents and better than the minimal set.
(a) Total causal change.
ParentsMinimalTheorem 1&2000.10.10.10.1Error
(b) Direct causal change.
ParentsMinimalTheorem 3&4000.10.10.10.1Error
Figure 3: Mean absolute error and standard deviation for estimating the total and direct causal changes of $X$ on $Y$. Subfigure (a) compares errors in total causal change estimation using (i) the set of parents of $X$, (ii) a minimal set satisfying the back-door criterion, and (iii) the set derived from Theorem 1 and 2 using a difference graph. Subfigure (b) evaluates direct causal change estimation using (i) the set of parents of $Y$, (ii) a minimal set satisfying the single-door criterion, and (iii) the set from Theorem 3 and 4. Results are reported over $100$ pairs of dataset, consisting of $1000$ observations and $10$ variables, with each $10$ pairs sharing the same difference graph for which the causal effect of interest is identifiable by our theorems. Each dataset is simulated using a linear SCM, where coefficients are randomly selected between 00 and $1$, with Gaussian noise.
5 Related works and discussion